Jamie has 300 ft of fencing to enclose a rectangular pasture. The pasture's length is to be 10 ft less than 3 times the width. Find the width of the garden.

A. Write a system of equations for the problem.
B. Solve your system. Make sure to show/explain your work for credit.
Please solve as quickly as possible

Answer by stanbon(75887)

Jamie has 300 ft of fencing to enclose a rectangular pasture. The pasture's length is to be 10 ft less than 3 times the width. Find the width of the garden.
-------------------------------
A. Write a system of equations for the problem.
Perimeter = 2(length + width)
300 = 2(3w-10 + w)
-------------------------
B. Solve your system.
150 = (4w-10
4w = 160
w = 40 ft (width)
Length = 3w-10 = 3*40-10 = 110 ft
------------------
Cheers,
Stan H.
=============